OpenStudy (anonymous):

What is the value of x in the solution to the following system of equations? x - 2y = 2 3x + y = 6

OpenStudy (anonymous):

Standard form: x-2y=2; 3x+y=6; substitute/eliminate x = +2y+2 substitute/eliminate y = -0 Solution: x = 2; y = 0; (x, y) = (2, 0)